OpenStudy (anonymous):

12.Given that f:x->3x-4,find the object of function that maps onto itself. @Conqueror

OpenStudy (anonymous):



OpenStudy (anonymous):

f(x)=3(2)-4 =6-4 =2

OpenStudy (anonymous):

Am i correct?

pooja195 (pooja195):

it looks correct to me but lets have @AnswerMyQuestions or @iGreen look over it

OpenStudy (anonymous):

or f(x)=3(-2)-4 =-6-4 =-10

OpenStudy (anonymous):

@Nishant_Garg

pooja195 (pooja195):

I would say the first one is right =/

OpenStudy (anonymous):

Thnx @pooja195

OpenStudy (anonymous):

\[y=f(x)=3x-4\] Given:f(x)=x\[x=3x-4\]\[2x=4\]\[x=2\]